i am supposed to identify like terms but i don't understand it my problem is -17-17=6-x you can see that the problem doesn't have like terms. I asked my teacher but she said that i still had to solve it. she told me though in class that it can have 1 answer many answers or no answer please help.
Found 2 solutions by rfer, ankor@dixie-net.com:
Answer by rfer(10417)

-17-17=6-x
all of the whole numbers are like terms
if you had more than one x they would be like terms
combine -17 and -17=-34
-34=6-x
subtract 6 from both sides
-34-6=6-6-x
-40=-x
we can't have a -x as part of the answer so multiply through by -1
40=x
x=40

Answer by ankor@dixie-net.com(12684)

-17 - 17 = 6 - x
which is
-34 = 6 - x
:
subtract 6 from both sides
-34 - 6 = -x
-40 = -x
:
x has to be positive, multiply both sides by -1 to change the signs
40 = x; the only answer
:
:
Check it in the original equation
- 17 - 17 = 6 - 40
-34 = -34, equality reigns!!